OpenStudy (anonymous):

5/9v+w=z, for v show work please?

OpenStudy (unklerhaukus):

\[ \frac5{9v}+w=z\]or\[ \frac5{9}v+w=z\]or\[ \frac5{9v+w}=z\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 9 }v+w=z\]

OpenStudy (unklerhaukus):

first take away w form both sides, then multiply both sides by 9/5